Rosie's Florist borrows $420,000 to be paid off in four years. The loan payments are semiannual with the first payment due in six months, and interest is at 6%. What is the amount of each payment?

Multiple Choice

  • $60,587.

  • $62,630.

  • $60,832.

  • $59,832.
Answer #1

The Answer is the last option $59,832

P Principal Amount = $ 420,000
No. of years = 4
n No. of instalmets = 4*2 = 8
( paid semi annually)
Interest rate per annum = 6% or 0.06
i Interest rate per period (half year) = (6/2)% = 3% or 0.03
PMT = Payment for period should be found
The formula for Present Value of Annuity is to be used
P = PMT x [ ( 1 – [ (1 + r) ^ -n ] ) / r ]
420,000 = PMT * [ ( 1 - [ (1+0.03) ^ -8] )/0.03 ]
420,000 = PMT * [ ( 1 - [ (1.03) ^ -8] )/0.03 ]
420,000 = PMT * [ ( 1 - [ 0.7894 ] )/0.03 ]
420,000 = PMT * [ ( 0.2106 )/0.03 ]
420,000 = PMT * [ 7.0197 ]
PMT = 420,000 / 7.0197
PMT = $ 59,832